OpenStudy (anonymous):

What is the value of (500,001) squared -(499,999) squared

OpenStudy (anonymous):

any tips for this problem?

hartnn (hartnn):

this is simple

OpenStudy (anonymous):

Difference of squares

hartnn (hartnn):

\( \huge a^2-b^2 = (a+b)(a-b)\)

OpenStudy (anonymous):

ahh

OpenStudy (anonymous):

so

OpenStudy (anonymous):

ehh

OpenStudy (anonymous):

ahh

OpenStudy (anonymous):

well

OpenStudy (anonymous):

if you subtract 500,001 and 499,999, you get 2

hartnn (hartnn):

yup

OpenStudy (anonymous):

what's the trick for this problem?

OpenStudy (anonymous):

add 6 zeros?

OpenStudy (anonymous):

cause 500,001 and 499,999=1,000,000

hartnn (hartnn):

what is 500001+499999=?

OpenStudy (anonymous):

1,000,000 times 2=2,000,000?

OpenStudy (anonymous):

is that the trick?

hartnn (hartnn):

u got it.

OpenStudy (anonymous):

find the difference of the 2 numbers and the sum then multiply?

hartnn (hartnn):

yes.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

next problem

OpenStudy (anonymous):

after tonks finishes writing

OpenStudy (anonymous):

500001^2 - 499999^2 = (500001 + 499999)(500001 - 499999) You already got it, I'm so slow!

OpenStudy (anonymous):

k

OpenStudy (anonymous):

unto the next problem :)

OpenStudy (anonymous):

these are PSAT ?s

OpenStudy (anonymous):

so you have to be fast